A projectile is fired at an 18 degree angle with an initial velocity of 11.3m/s. How far fro its launch will it land assuming the ground is level?
1
Expert's answer
2020-10-12T07:48:45-0400

The formula for the range of projectile:


"d=\frac{v^2}{g}\sin{2\theta}\\d=\frac{11.3^2}{9.8}\sin{36}=7.66\ m"